Apple's market share in the smart phone market represents a proportion of 0.1935. If Apple conducts a survey nationwide of 522 smart phone users, what is the probability that no more than 93 of the people are Apple users?
Answer #1

Solution :

Using binomial distribution,

\mu = n * p = 0.1935 * 522 = 101.007

\sigma = \sqrt n * p * q = \sqrt 0.8065 * 0.1935 * 522 =  9.02564

Using continuity correction ,

= P((x - \mu ) / \sigma < (93.5 - 101.007) / 9.02564)

= P(z < -0.83)

= 0.2033

Probability = 0.2033